OpenStudy (anonymous):

Salisbury High School decided to take their students on a field trip to a theme park. A total of 150 people went on the trip. Adults pay $45.00 for a ticket and students pay $28.50 for a ticket. How many students and how many adults went to the park if they paid a total of $4770? And so far I have: A= # of adults C= # of students A+C= 150 A45+C28.5=4770 But I'm having trouble using these equations to find the values of A and C. I think that I might divide 4770 by either 28.5 or 45, but I'm not sure where I would go from there if I was right.

OpenStudy (campbell_st):

well is you want to use elimination, them multiply the 1st equation by 45 and then subtract the 2nd equation. This will eliminate A and allow you to solve for C 45A + 45C = 6750 - 45A +28.5C = 4770 ---------------------- 16.5 C = 1980 now solve for C substitute that value into the 1st equation to find A hope it helps